Operation Summary for October 30, 2009 - November 2, 2009 |
The Integrated Luminosity for the period from 10/26/09 to 11/2/09 was 45.48 pb-1. Friday Oct. 30 The day shift began with Operations monitoring the Tevatron (TeV) store 7306 (60.74E30) and the Recycler stash (170.9E10), with the Antiproton Source stacking (17.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations reported that the MI-52 kicker current was unsteady due to humidity; MI personnel investigated. Operations and MTA personnel performed MTA MuCool experiment beam permit and power supply tests. TeV personnel began an end-of-store beam study at the end of the shift. During the evening shift, TeV personnel conducted crystal collimator studies. Problems with a MI RF station (MIRF2) caused NuMI beam trips; operators adjusted the MIRF2 feed-forward system to resolve the problem. A Lambertson (LAM61) tripped off on an interlock indication; power supply technicians came in and adjusted the power supply’s door interlock switches. Operations put the TeV into shot setup and established store 7307 at 9:04 PM with an average initial luminosity (AIL) of 279.52E30. Saturday Oct. 31 The midnight shift began with Operations monitoring store 7307 (166.03E30) and the stash (122.5E10), with Pbar stacking (13.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, the Pelletron developed a problem. Recycler personnel investigated and reported that the Pelletron would need to be opened for repairs. During the day shift, at 11 AM, the H- Source’s output dropped; operators conferred with experts and then raised its pressure, which restored its output. Store 7307 aborted when a TeV sector A2 small transformer blew its fuses and caused the VCB to open; TeV personnel investigated and made repairs. CDF and D0 access their collision halls. Mechanical techs emptied the Pelletron of its insulating gas in preparation for an access and repairs. During the evening shift, Operations put the TeV through a dry squeeze and then into shot setup. At 8:03 PM the TeV suffered at quench at sector A1. Operators and TeV personnel investigated. Recycler personnel accessed MI-31 and began Pelletron repairs. Sunday Nov. 1 The midnight shift began with TeV personnel investigating the quench, with Operations monitoring the stash, with Pbar stacking, with Recycler personnel making Pelletron repairs, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TeV into shot setup and established store 7315 at 2:46 AM with an AIL of 74.69E30. Operations reported that the MI vacuum (MIVAC) front end needed three reboots. Zombie gore cleaned up from failed invasion attempt. During the day shift, Operations terminated store 7315 at 1 PM and put the TeV into shot setup. Operations established store 7316 at 2:45 PM with and AIL of 84.5E30. LRF3 repeatedly suffered from driver and power amplifier trips. During the evening shift, the Pelletron work continued. LRF3 trips continued; Linac personnel instructed operators on how to resolve the problems. Operations continued to reboot the MIVAC front end. Mechanical techs reported finding a leak on a Pelletron flange. Since this work will add another 5-6 hours to the job, the techs went home for the night; they will resume work on Monday morning. Monday Nov. 2 The midnight shift began with Operations monitoring store 7316 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7316 at 2:02 AM. Operations put the TeV into shot setup at established store 7317 at 3:27 AM with an AIL of 85.57E30. Operations had to reboot the MIVAC once. Operations reported a problem with BRF4 due to a power amplifier LCW leak indication; Booster personnel investigated. The Plans for Monday The plans for today are to stack, stash, and store, and to get the Pelletron back online by this afternoon. The Run Coordinator will allow Booster and MI personnel to conduct studies and for MTA personnel to conduct further power supply and beam permit tests. |
|
More Information Comments and Suggestions |
| last modified 11/02/2009 email Fermilab |
| Security, Privacy, Legal | |